# cuVSLAM: CUDA accelerated visual odometry and mapping

### v2

@@ 2 System Architecture and Design
In contrast to the frontend, cuVSLAM’s backend focuses on global map and pose consistency. It processes frontend outputs and builds a globally consistent map that integrates camera poses, 2D observations, 3D landmarks, pose deltas, and visual features. The backend maintains global consistency in the map through asynchronous refinement using pose graph optimization and loop closing.
@@ 2.2.2 Multi-Stereo
The FIG is built automatically at cuVSLAM’s startup. Provided with a list of camera extrinsics/intrinsics, cuVSLAM iterates over each pair of cameras and creates a new edge in the graph when cameras have overlapping fields of view (FoV). To detect this, cuVSLAM places a virtual plane in front of the first camera. It selects uniformly distributed pixels in the image $I_{1}$, lifts them onto the plane, and backprojects them to the image $I_{2}$. When the ratio of successfully projected points to the total number of sampled points exceeds a threshold, cuVSLAM detects the shared FoV and creates an edge.
@@ 2.3 Loop closing and global map refinement
cuVSLAM’s loop closing and global map refinement module helps limit global drift by enforcing global consistency of the map. It consists of 6 DoF keyframe poses, landmarks and their observations, pose deltas, and sparse image features. Pose deltas are obtained through the integration of intra-keyframe frontend pose predictions. For image features, we use a list of $9\times 9$ image patches taken from each level of the image pyramid.
@@ 3.1 System Performance and Hardware Utilization
We evaluated cuVSLAM’s performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware. The test scenarios involved processing feature-rich images captured from either a wheeled robot navigating a warehouse environment or handheld camera movements. To ensure accurate measurements, all experiments were conducted on systems dedicated solely to cuVSLAM execution, without concurrent resource-intensive operations such as visualization, complex image processing, or data recording that could potentially skew the results.
Table 1 summarizes the key performance metrics, with comprehensive results provided in Appendix A.2. The left portion of Table 1 presents execution times for cuVSLAM’s visual tracking function, excluding image acquisition and post-processing overhead. These measurements were obtained using recorded color images at 768×480 resolution (with the exception of mono-depth mode at 640×480 resolution) on two distinct hardware configurations: a desktop workstation featuring an Intel i7-14700 CPU with NVIDIA RTX 4090 GPU, and an embedded Jetson Orin AGX 64GB single-board computer operating in MAXN power mode.
@@
The right portion of Table 1 details CPU and GPU utilization percentages for cuVSLAM visual odometry executed through ISAAC ROS 3.2 with RealSense D435/D455 stereo cameras. All live tests utilized cameras configured at 640×480 resolution and 60 FPS, except for Stereo-Inertial mode, which ran at 30 FPS with IMU data at 200 Hz. The reduced framerate was necessary to ensure valid IMU-based integration, providing 6-7 IMU measurements between each stereo image pair. To accurately isolate cuVSLAM’s hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active. Resource utilization metrics were sampled at 15 ms intervals for equal durations in both experiments. The net resource consumption attributable to cuVSLAM was calculated by subtracting baseline measurements from total utilization values and averaging the results. This methodology provides an accurate assessment of the computational overhead introduced by the library. Additional evaluation details, including performance benchmarks for cuVSLAM executed via Python API on entry-level Jetson Orin Nano hardware, are presented in Appendix A.2.
@@ 3.2 Benchmarking
We evaluated cuVSLAM’s different operation modes using several publicly available datasets. The Mono-Depth mode was validated on the ICL-NUIM Handa et al. (2014) and TUM RGB-D Sturm et al. (2012) datasets. The Stereo mode was tested on the EuRoC Burri et al. (2016) and KITTI Geiger et al. (2013) datasets. The Stereo-Inertial mode was assessed using the EuRoC dataset. The Multi-Stereo mode was evaluated on simulated TartanAir V2 Wang et al. (2023) and TartanGround Patel et al. (2025) datasets and our proprietary R2B real-world dataset.
To quantify trajectory estimation accuracy, we employed standard metrics: Average Relative Translation and Rotation Errors Geiger et al. (2013), and Absolute Pose Error with trajectory alignment and scale correction Sturm et al. (2012) using the EVO library Grupp (2017).
@@
Mode Dataset Method ATE, % ARE, deg APE (RMSE)
Mono-Depth ICL-Nuim ORBSLAM3 - - -
@@
To ensure accurate and representative comparisons, we excluded clearly identified outlier and corrupted sequences. For instance, the EuRoC V203 sequence was included only in Stereo-Inertial evaluations, and only a subset of TUM RGB-D sequences were used (see Table 4). For TartanGround datasets, not all sequences provide the full complement of cameras or are available for download. Additionally, many sequences lack overlapping trajectories, preventing loop closure detection—a critical component for evaluating SLAM performance versus odometry. Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 7) and TartanAir V2 (see Tables 9 and 9) datasets.
@@
While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, we needed to rigorously validate cuVSLAM’s multi-camera capabilities in real-world scenarios. We therefore collected a proprietary Multi-Stereo R2B dataset in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras. The example views from the R2B dataset are shown in Fig. 7. This was our only available option for validating this cuVSLAM mode on real data, as publicly available real-world multi-camera datasets either lack hardware-synchronized cameras suitable for stereo pair arrangements Carlevaris-Bianco et al. (2015); Nair et al. (2024) or employ software synchronization that cannot ensure simultaneous image capture Yin et al. (2021); Zhang et al. (2023).
Table 3 presents cuVSLAM’s results in Multi-Stereo Camera Mode. Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results. For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities. Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 6. While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM
Dataset Method ATE, % ARE, deg APE (RMSE)
@@ 4.1 Trajectory Reliability Under Visual Occlusion
Figure 8 demonstrates the system’s trajectory reliability through a practical robustness test. We equipped a Carter robot with 4 stereo cameras and subjected it to challenging visual conditions during continuous motion. Cameras were randomly occluded with opaque film for intervals of 20–60 seconds, with the constraint that at least one stereo pair remained unobstructed at a given moment. Despite these severe visual constraints, the odometry trajectory maintained stability without exhibiting jumps or discontinuities, validating the system’s robustness to partial visual obstruction. This capability represents a significant advancement for visual tracking systems. While single-camera configurations can maintain tracking in feature-poor environments only briefly—typically relying on IMU or depth sensor fusion—our multi-camera approach preserves stable and reliable visual tracking over extended periods even with degraded visual input.
Figure 8: Occlusion resilience test results using the Carter robot. The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film. Numbered image rows correspond to instantaneous views from the robot’s 4 stereo cameras (showing only the left image from each stereo pair) at the numbered locations along the trajectory.
@@ A.1.1 KITTI Dataset
We evaluated our method on sequences 0-10 of the KITTI dataset, for which ground truth trajectories are available. For comparison with the conventional KITTI leaderboard (1), Figure 10 presents relative translation and rotation errors calculated on trajectory segments of lengths 100, 200, 300, 400, 500, 600, 700, and 800 meters. In contrast Table 2 presents results calculated without trajectory segmentation to maintain metric comparability across different datasets.
Figure 10: cuVSLAM evaluation results on KITTI odometry benchmark sequences 00-10. Translation and rotation errors are calculated on segments of 100-800m following the KITTI public leaderboard methodology, averaged metrics: ATE: 0.85%, ARE: 0.0025 [deg/m].
A.1.2 TUM RGB-D Dataset
To validate the Mono-Depth tracking mode, we employed a subset of 10 sequences from the Freiburg3 collection of the TUM RGB-D dataset (20). These sequences were selected because they provide rectified images, which are required by the cuVSLAM Mono-Depth mode. We performed minimal preprocessing, implementing only depth-image pair synchronization without any additional image or depth processing. The complete list of evaluated sequences is presented in Table 4.
@@ A.1.2 TUM RGB-D Dataset
Sequence Mode ATE ARE APE
@@
A.1.3 TartanGround and TartanAir V2
Odometry Mode. We validated cuVSLAM using 216 sequences from TartanGround  (18) and 539 sequences from TartanAir V2  (21) Hard datasets. Complete sequence lists are provided in Tables 5 and 9, respectively. All results were obtained using a four stereo-camera configuration (front, back, left, and right) with 640×640 undistorted images. While the TartanAir V2 dataset supports two additional cameras (top and bottom), their inclusion yielded minimal metric improvement (less than 5%). For the TartanGround dataset, we observed slight performance degradation when using these additional cameras, attributed to meaningless feature selection from sky regions and repetitive floor textures in planar robot navigation areas.
To facilitate comparison with other visual tracking libraries, we calculated additional metrics—relative translation and rotation per frame—using the official TartanAir GitHub repository (2). As not all environments mentioned in the original TartanGround paper were available for download, we provide environment-averaged results to enable future comparative studies.
@@ A.1.3 TartanGround and TartanAir V2
Environment Sequences ATE ARE APE $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
@@
Environment Difficulty ATE ARE APE
@@
SLAM Mode. To evaluate the impact of SLAM mode on trajectory estimation, we analyzed only sequences containing loop closures. This resulted in a subset of 32 sequences for TartanGround and 506 sequences for TartanAir V2 Hard. Complete tracking results are provided in Tables 7 and 9, with all results averaged across environments.
@@
Environment Sequences Mode ATE ARE APE
@@
A.1.4 Mutli-Stereo R2B Dataset
Figure 12: Trajectories predicted by cuVSLAM (4 Multi-Stereo Odometry) on the proprietary R2B wheeled robot dataset. Black lines represent ground truth; green lines represent odometry predictions. Attached table lists trajectory distances, number of 8-image batches, and trajectory numbering per environment (left to right).
A.2 Appendix B: Hardware Scalability on Embedded Platforms
Jetson Orin Nano Super Performance. To evaluate the feasibility of deploying cuVSLAM on entry-level hardware, we conducted experiments replacing the Jetson Orin AGX with a Jetson Orin Nano Super operating in 25W mode and utilizing exclusively the Python stack. The experiments explore possibility of using multi-stereo odometry mode for 640x360 resolution and 30 FPS followed the methodology described above with free camera movement.
Table 8: TartanAir V2 cuVSLAM evaluation results (Envs 1–35, see Appendix A.1.3 for details)
Env Seq Mode ATE ARE APE
@@ A.2 Appendix B: Hardware Scalability on Embedded Platforms
Table 9: TartanAir V2 cuVSLAM evaluation results (Envs 36–71, see Appendix A.1.3 for details)
Env Seq Mode ATE ARE APE

### v3

@@ 2 System Architecture and Design
In contrast to the frontend, cuVSLAM backend focuses on global map and pose consistency. It processes frontend outputs and builds a globally consistent map that integrates camera poses, 2D observations, 3D landmarks, pose deltas, and visual features. The backend maintains global consistency in the map through asynchronous refinement using pose graph optimization and loop closing.
@@ 2.2.2 Multi-Stereo
The FIG is built automatically at cuVSLAM startup. Provided with a list of camera extrinsics/intrinsics, cuVSLAM iterates over each pair of cameras and creates a new edge in the graph when cameras have overlapping fields of view (FoV). To detect this, cuVSLAM places a virtual plane in front of the first camera. It selects uniformly distributed pixels in the image $I_{1}$, lifts them onto the plane, and backprojects them to the image $I_{2}$. When the ratio of successfully projected points to the total number of sampled points exceeds a threshold, cuVSLAM detects the shared FoV and creates an edge.
@@ 2.3 Loop closing and global map refinement
cuVSLAM loop closing and global map refinement module helps limit global drift by enforcing global consistency of the map. It consists of 6 DoF keyframe poses, landmarks and their observations, pose deltas, and sparse image features. Pose deltas are obtained through the integration of intra-keyframe frontend pose predictions. For image features, we use a list of $9\times 9$ image patches taken from each level of the image pyramid.
@@ 3.1 System Performance and Hardware Utilization
We evaluated cuVSLAM performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware. The test scenarios involved processing feature-rich images captured from either a wheeled robot navigating a warehouse environment or handheld camera movements. To ensure accurate measurements, all experiments were conducted on systems dedicated solely to cuVSLAM execution, without concurrent resource-intensive operations such as visualization, complex image processing, or data recording that could potentially skew the results.
Table 1 summarizes the key performance metrics, with comprehensive results provided in Appendix A.2. The left portion of Table 1 presents execution times for cuVSLAM visual tracking function, excluding image acquisition and post-processing overhead. These measurements were obtained using recorded color images at 768×480 resolution (with the exception of mono-depth mode at 640×480 resolution) on two distinct hardware configurations: a desktop workstation featuring an Intel i7-14700 CPU with NVIDIA RTX 4090 GPU, and an embedded Jetson Orin AGX 64GB single-board computer operating in MAXN power mode.
@@
The right portion of Table 1 details CPU and GPU utilization percentages for cuVSLAM visual odometry executed through ISAAC ROS 3.2 with RealSense D435/D455 stereo cameras. All live tests utilized cameras configured at 640×480 resolution and 60 FPS, except for Stereo-Inertial mode, which ran at 30 FPS with IMU data at 200 Hz. The reduced framerate was necessary to ensure valid IMU-based integration, providing 6-7 IMU measurements between each stereo image pair. To accurately isolate cuVSLAM hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active. Resource utilization metrics were sampled at 15 ms intervals for equal durations in both experiments. The net resource consumption attributable to cuVSLAM was calculated by subtracting baseline measurements from total utilization values and averaging the results. This methodology provides an accurate assessment of the computational overhead introduced by the library. Additional details on evaluation and an exploration of hardware utilization across different resolutions and modes on Jetson devices are provided in Appendix A.2.
@@ 3.2 Benchmarking
We evaluated cuVSLAM different operation modes using several publicly available datasets. The Mono-Depth mode was validated on the AR-table Chen et al. (2023), ICL-NUIM Handa et al. (2014) and TUM RGB-D Sturm et al. (2012) datasets. The Stereo mode was tested on the EuRoC Burri et al. (2016) and KITTI Geiger et al. (2013) datasets. The Stereo-Inertial mode was assessed using the EuRoC Burri et al. (2016) and TUM-VI Schubert et al. (2018) datasets. The Multi-Stereo mode was evaluated on simulated TartanAir V2 Wang et al. (2023) and TartanGround Patel et al. (2025) datasets and our proprietary R2B real-world dataset.
@@
Mode Dataset Method avgRTE, % avgRE, deg RMSE APE
Mono-Depth AR table ORBSLAM3 - - -
DPVO 4.42 1.65 0.77
cuVSLAM Odom 0.34 3.59 0.09
cuVSLAM SLAM 0.19 1.68 0.025
ICL-Nuim ORBSLAM3 - - -
@@
TUM-VI Room ORBSLAM3 2.17 1.37 0.077
DPVO - - -
cuVSLAM Odom 0.20 3.85 0.18
cuVSLAM SLAM 0.12 3.00 0.12
To quantify visual tracking accuracy, we employed standard metrics: Average Relative Translation (avgRTE) and Rotation (avgRE) Errors Geiger et al. (2013), and Absolute Pose Error (RMSE APE) with trajectory alignment and scale correction Sturm et al. (2012) using the EVO library Grupp (2017).
@@
To ensure accurate and representative comparisons, we excluded clearly identified outlier and corrupted sequences. For instance, the EuRoC V203 sequence was included only in Stereo-Inertial evaluations, and only a subset of TUM RGB-D sequences were used (see Table 6). For TartanGround datasets, not all sequences provide the full complement of cameras or are available for download. Additionally, many sequences lack overlapping trajectories, preventing loop closure detection—a critical component for evaluating SLAM performance versus odometry. Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 9) and TartanAir V2 (see Tables 11 and 11) datasets.
While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, rigorous validation of cuVSLAM multi-camera capabilities in real-world scenarios was required. Therefore, a proprietary Multi-Stereo R2B dataset was collected in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras. The example views from the R2B dataset are shown in Fig. 7. This was our only available option for validating this cuVSLAM mode on real data, as publicly available real-world multi-camera datasets either lack hardware-synchronized cameras suitable for stereo pair arrangements Carlevaris-Bianco et al. (2015); Nair et al. (2024) or employ software synchronization that cannot ensure simultaneous image capture Yin et al. (2021); Zhang (2023).
Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM on multi-camera datasets
Dataset Method avgRTE, % avgRE, deg RMSE APE
@@
Table 3 presents cuVSLAM results in Multi-Stereo Mode. Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results. For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities. Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 8. While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
@@ 4.1 Trajectory Reliability Under Visual Occlusion
Figure 8 demonstrates the system’s trajectory reliability through a practical robustness test. We equipped a Carter robot with 4 stereo cameras and subjected it to challenging visual conditions during continuous motion. Cameras were randomly occluded with opaque film for intervals of 20–60 seconds, with the constraint that at least one stereo pair remained unobstructed at a given moment.
Figure 8: Occlusion resilience test using the Carter robot. cuVSLAM maintains continuous trajectory despite random camera occlusions. Numbered rows show instantaneous views from the robot’s 4 stereo cameras (left images only) at corresponding trajectory locations
Despite these severe visual constraints, the odometry trajectory maintained stability without exhibiting jumps or discontinuities, validating the system’s robustness to partial visual obstruction. This capability represents a significant advancement for visual tracking systems. While single-camera configurations can maintain tracking in feature-poor environments only briefly—typically relying on IMU or depth sensor fusion—our multi-camera approach preserves stable and reliable visual tracking over extended periods even with degraded visual input.
@@ A.1.1 KITTI Dataset
We evaluated our method on sequences 0-10 of the KITTI dataset, for which ground truth trajectories are available. For comparison with the conventional KITTI leaderboard (1), Figure 10 presents relative translation (avgRTE) and rotation (avgRE) errors calculated on trajectory segments of lengths 100, 200, 300, 400, 500, 600, 700, and 800 meters. In contrast Table 2 presents results calculated without trajectory segmentation to maintain metric comparability across different datasets.
Figure 10: cuVSLAM evaluation results on KITTI odometry benchmark sequences 00-10. Translation and rotation errors are calculated on segments of 100-800m following the KITTI public leaderboard methodology, averaged metrics: Translation: 0.85%, Rotation: 0.0025 [deg/m].
A.1.2 TUM-VI Room Dataset
To validate the stereo-inertial tracking mode on real hardware under aggressive camera motions, we evaluated cuVSLAM on the Room sequences of the TUM-VI dataset (20). These sequences were selected because they provide complete ground truth trajectories. We used undistorted images with a resolution of 512x512 pixels and applied a 50-pixel mask along each frame edge to mitigate significant fisheye lens distortion, which negatively affects tracking performance. To facilitate comparison with other visual tracking methods, we followed the evaluation protocol described in the original TUM-VI paper (20), computing the average RMSE of the Relative Pose Error (RPE) in meters/degrees over 1-second segments. Detailed results are provided in Table 4.
Table 4: cuVSLAM Stereo Inertial tracking validation results on TUM-VI Room dataset RMSE RPE (m / deg) on 1 second segments
Sequence ODOM SLAM
room1 0.023 / 0.89 0.025 / 0.90
room2 0.048 / 2.28 0.048 / 2.29
room3 0.047 / 1.93 0.048 / 1.95
room4 0.015 / 0.54 0.015 / 0.55
room5 0.019 / 0.56 0.021 / 0.57
room6 0.021 / 0.57 0.021 / 0.58
A.1.3 AR-table Dataset
This dataset contains sequences captured using an Intel RealSense D455 depth camera, providing image and depth data at a resolution of 848x480 pixels. The enhanced image and depth quality allows achieving performance on this real-world camera close to that obtained on the ideal simulated ICL-NUIM dataset, as shown in Table 2.
However, consistent gaps in depth values were observed along the sides and bottom edges of the depth images. To mitigate this issue, we applied static masking of 24 pixels on each side and 10 pixels at the bottom. Additionally, several prolonged depth-image dropouts (lasting more than one second) occurred in sequences 3, 4, 5, and 8. To address these interruptions, we segmented the affected sequences into continuous portions, executed tracking independently on each segment, and subsequently merged the results.
To facilitate comparison with other visual tracking methods, we followed the evaluation methodology described in the original AR-table dataset paper (8). Specifically, we computed Absolute Trajectory Error (ATE), reporting translation errors in centimeters and rotation errors in degrees. The obtained results are summarized in Table 5.
Table 5: cuVSLAM Mono-Depth Odometry and SLAM validation results on AR table dataset ATE (degree / cm)
Mode table_01 table_02 table_03 table_04 table_05 table_06 table_07 table_08
ODOM 3.77 / 2.03 9.36 / 5.41 6.09 / 2.28 3.13 / 2.21 2.27 / 1.97 3.43 / 3.39 18.5 / 6.17 13.2 / 8.18
SLAM 2.98 / 1.76 2.23 / 1.77 2.11 / 0.87 2.12 / 1.64 1.24 / 1.22 1.93 / 1.23 8.88 / 3.82 4.83 / 3.69
A.1.4 TUM RGB-D Dataset
To validate the Mono-Depth mode on this dataset, a subset of 10 sequences from the Freiburg3 collection was used. These sequences were selected as they provide rectified images, which are required by cuVSLAM Mono-Depth mode. Minimal preprocessing was performed, consisting only of depth-image pair synchronization without further image or depth processing. The complete list of evaluated sequences is provided in Table 6.
@@ A.1.4 TUM RGB-D Dataset
Sequence Mode avgRTE avgRE RMSE APE
@@
A.1.5 TartanGround and TartanAir V2
Odometry Mode. We validated cuVSLAM using 216 sequences from TartanGround  (19) and 539 sequences from TartanAir V2  (23) Hard datasets. Complete sequence lists are provided in Tables 7 and 11, respectively. All results were obtained using a four stereo-camera configuration (front, back, left, and right) with 640×640 undistorted images. While the TartanAir V2 dataset supports two additional cameras (top and bottom), their inclusion yielded minimal metric improvement (less than 5%). For the TartanGround dataset, we observed slight performance degradation when using these additional cameras, attributed to meaningless feature selection from sky regions and repetitive floor textures in planar robot navigation areas.
To facilitate comparison with other visual tracking libraries, we computed additional metrics—relative translation ($t_{\mathrm{rel}}$) and rotation ($r_{\mathrm{rel}}$) per frame—using scripts from the official TartanAir GitHub repository (2). Since not all environments from the original TartanGround paper were available, we present environment-averaged results in Table 7 to enable future comparative studies.
@@ A.1.5 TartanGround and TartanAir V2
Environment Sequences avgRTE avgRE RMSE APE $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
@@
Environment Difficulty avgRTE avgRE RMSE APE
@@
SLAM Mode. To evaluate the impact of SLAM mode on trajectory estimation, we analyzed only sequences containing loop closures. This resulted in a subset of 32 sequences for TartanGround and 506 sequences for TartanAir V2 Hard. Complete tracking results are provided in Tables 9 and 11, with all results averaged across environments.
A.1.6 Mutli-Stereo R2B Dataset
Figure 12: Trajectories predicted by cuVSLAM (4 Multi-Stereo Odometry) on the proprietary R2B wheeled robot dataset. Black lines represent ground truth; green lines represent odometry predictions. Attached table lists trajectory distances, number of 8-image batches, and trajectory index for environments (left to right).
A.2 Appendix B: Hardware Scalability on Embedded Platforms
Jetson Orin Nano Performance. To evaluate the feasibility of deploying cuVSLAM on entry-level hardware, we conducted experiments replacing the Jetson Orin AGX with a Jetson Orin Nano Super operating in 25W mode and utilizing exclusively the Python stack. The experiments explore possibility of using multi-stereo odometry mode for 640x360 resolution and 30 FPS followed the methodology described above with free camera movement.
@@ A.2 Appendix B: Hardware Scalability on Embedded Platforms
Environment Sequences Mode avgRTE avgRE RMSE APE
@@
Table 10: TartanAir V2 Hard cuVSLAM evaluation results (Envs 1–35, see Appendix A.1.5 for details)
Env Seq Mode RTE RRE APE
@@
Table 11: TartanAir V2 Hard cuVSLAM evaluation results (Envs 36–71, see Appendix A.1.5 for details)
Env Seq Mode RTE RRE APE